# Forbidding just one intersection for short integer sequences

### v1

@@ 4 Cross-disagreement in homogeneous families
Now we are ready to proof a measure boosting lemma. The lemma is a counterpart of Lemma $6.11$ from [keevash2023forbidden].
@@
The formal proof of Lemma 27 is quite technical, therefore, we give its sketch first. We make several modifications of families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ to get families with much bigger measure on the same groud set $m^{\prime}]^{n^{\prime}}$ with $m^{\prime}\ll m$, such that if the modified families contain cross-disagreement than the initial families also contain cross-disagreement. Then we use Lemma 11 to find cross-disagreement in the modified families. The proof is divided into the following steps.
@@
Step $4$. We repeat steps 11-33, switching families. That is, we boost the measure of $\mathcal{F}^{\prime\prime}_{2}$ and ensure that the measure of $\mathcal{F}^{\prime}_{1}$ decreases insignificantly. After this step, we have families $\mathcal{F^{\prime\prime\prime}}_{1}\subset[m^{\prime\prime}]^{[n]\setminus(H_{1}\cup R\cup R^{\prime})}$ and $\mathcal{F^{\prime\prime\prime}}_{2}\subset[m^{\prime\prime}]^{[n]\setminus(H_{2}\cup R\cup R^{\prime})}$ with constant measures such that the existence of cross-disagreement in $\mathcal{F^{\prime\prime\prime}}_{1}$ and $\mathcal{F^{\prime\prime\prime}}_{1}$ implies existence of cross-disagreement in the initial families.
@@
$|R^{\prime}|<\frac{\ln\left(2\alpha_{2}^{-1}\right)}{\ln 2}$, and $(\nu^{\prime})^{\pi^{\prime}_{-H_{2}}}_{R^{\prime}\rightarrow y^{\prime}}(\mathcal{F^{\prime\prime}}_{2}^{\pi^{\prime}_{-H_{2}}}(R^{\prime}\rightarrow y^{\prime}))\geqslant\frac{1}{2}$. Let $\pi^{\prime}\in\Pi_{m,m^{\prime},b}^{\otimes[n]\setminus R}$ be arbitrary such that $\pi^{\prime}_{[n]\setminus(R\cup H_{2})}=\pi^{\prime}_{-H_{2}}$. Let $\mathcal{F^{\prime\prime\prime}}_{2}=\mathcal{F^{\prime\prime}}_{2}^{\pi^{\prime}}(R^{\prime}\rightarrow y^{\prime})$ and $\nu^{\prime\prime}=\nu^{\prime\pi^{\prime}}_{R^{\prime}\rightarrow y^{\prime}}=\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup R^{\prime})}$.
@@
Take arbitrary $x^{\prime}\in[m^{\prime\prime}]^{R^{\prime}\setminus H_{1}}$ such that $\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup R^{\prime}\cup H_{1})}(\mathcal{F^{\prime\prime}}_{1}^{\pi^{\prime}}(R^{\prime}\setminus H_{1}\rightarrow x^{\prime}))\geqslant\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup H_{1})}(\mathcal{F^{\prime\prime}}_{1}^{\pi^{\prime}})$. Note, that such $x^{\prime}$ exists by Claim 16 and $\operatorname{agr}(x^{\prime},y^{\prime})=0$, because for all $x^{\prime}$ with $\operatorname{agr}(x^{\prime},y^{\prime})\neq 0$ we have $\mathcal{F^{\prime\prime}}_{1}^{\pi^{\prime}}(R^{\prime}\setminus H_{1}\rightarrow x^{\prime})=\emptyset$. Let $\mathcal{F^{\prime\prime\prime}}_{1}=\mathcal{F^{\prime\prime}}_{1}^{\pi^{\prime}}(R^{\prime}\setminus H_{1}\rightarrow x^{\prime})$. By Claim 5 we get
$\nu^{\prime\prime}_{[n]\setminus(R\cup R^{\prime}\cup H_{1})}(\mathcal{F^{\prime\prime\prime}}_{1})=\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup R^{\prime}\cup H_{1})}(\mathcal{F^{\prime\prime\prime}}_{1})\geqslant\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup H_{1})}(\mathcal{F^{\prime\prime}}_{1}^{\pi^{\prime}})\geqslant\nu^{\prime}_{[n]\setminus(R\cup H_{1})}(\mathcal{F^{\prime\prime}}_{1})\geqslant\frac{1}{4}.$
@@ 5 Case $m\geqslant\operatorname{poly}(t)$, $n\geqslant\operatorname{poly}(t)\log m$
We need to check that conditions of Corollary 32 imply the conditions of Theorem 28. The case $t=1$, that is, the case of intersecting family of codes, is considered in [EKR_for_codes]. Thus, we may assume $t\geqslant 2$. Requirements (6), (7), (8) from Theorem 28 have a form $m\geqslant\text{poly}(t,\log m)$ and the polynomial from the requirement (7) has the highest degree $2C_{2}(2)<300$ in $t$. Thus, there exists a polynomial $m_{0}(t)$ of degree at most $300$ such that $m\geqslant m_{0}(t)$ implies requirements (6), (7), (8). Finally, for $m>e$ the condition (5) is a consequence of a condition $n>n_{0}(t)\ln(m)$, where
@@ Appendix A Proof of Lemma 12
Lemma 12 is a quantitative analogue of Theorem 2.3 from [keevash2023forbidden] and may be obtained by a verbatim repetition of its proof. We will prove the lemma by contradiction in the following way:
@@
One part of the transformation in Step 1 will be the following compression operator, defined in Section 4.2 of [keevash2023forbidden].
@@
Definition 9 (Definition 4.2 from [keevash2023forbidden]).
@@
The transformation at Step 1 will have the following properties, that are direct consequences of Fact 4.1 and Fact 4.3 from [keevash2023forbidden]
@@
For Step 2, we will need the following lemma of Ellis, Keller, and Lifshitz [ellis2019stability] to obtain a lower bound on $\mu_{1/2}$ from $\mu_{1/m}$.
@@
Lemma 46 (Fact 4.11 from [keevash2023forbidden]).

### v2

@@ 4 Cross-disagreement in homogeneous families
Now we are ready to proof a measure boosting lemma. The lemma is a counterpart of Lemma $6.11$ from [42].
@@
The formal proof of Lemma 27 is quite technical; therefore, we give its sketch first. We make several modifications of families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ to get families with much bigger measure on the same groud set $m^{\prime}]^{n^{\prime}}$ with $m^{\prime}\ll m$, such that if the modified families contain cross-disagreement than the initial families also contain cross-disagreement. Then we use Lemma 11 to find cross-disagreement in the modified families. The proof is divided into the following steps.
@@
Step $4$. We repeat steps 11-33, switching families. That is, we boost the measure of $\mathcal{F}^{\prime\prime}_{2}$ and ensure that the measure of $\mathcal{F}^{\prime}_{1}$ decreases insignificantly. After this step, we have families $\mathcal{F}^{\prime\prime\prime}_{1}\subset[m^{\prime\prime}]^{[n]\setminus(H_{1}\cup R\cup R^{\prime})}$ and $\mathcal{F}^{\prime\prime\prime}_{2}\subset[m^{\prime\prime}]^{[n]\setminus(H_{2}\cup R\cup R^{\prime})}$ with constant measures such that the existence of cross-disagreement in $\mathcal{F}^{\prime\prime\prime}_{1}$ and $\mathcal{F}^{\prime\prime\prime}_{1}$ implies the existence of cross-disagreement in the initial families.
@@
$|R^{\prime}|<\frac{\ln\left(2\alpha_{2}^{-1}\right)}{\ln 2}$, and $(\nu^{\prime})^{\pi^{\prime}_{-H_{2}}}_{R^{\prime}\rightarrow y^{\prime}}({\mathcal{F}_{2}^{\prime\prime}}^{\pi^{\prime}_{-H_{2}}}(R^{\prime}\rightarrow y^{\prime}))\geqslant\frac{1}{2}$. Let $\pi^{\prime}\in\Pi_{m,m^{\prime},b}^{\otimes[n]\setminus R}$ be arbitrary such that $\pi^{\prime}_{[n]\setminus(R\cup H_{2})}=\pi^{\prime}_{-H_{2}}$. Let $\mathcal{F}^{\prime\prime\prime}_{2}={\mathcal{F}_{2}^{\prime\prime}}^{\pi^{\prime}}(R^{\prime}\rightarrow y^{\prime})$ and $\nu^{\prime\prime}=\nu^{\prime\pi^{\prime}}_{R^{\prime}\rightarrow y^{\prime}}=\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup R^{\prime})}$.
@@
Take arbitrary $x^{\prime}\in[m^{\prime\prime}]^{R^{\prime}\setminus H_{1}}$ such that $\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup R^{\prime}\cup H_{1})}({\mathcal{F}^{\prime\prime}_{1}}^{\pi^{\prime}}(R^{\prime}\setminus H_{1}\rightarrow x^{\prime}))\geqslant\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup H_{1})}({\mathcal{F}_{1}^{\prime\prime}}^{\pi^{\prime}})$. Note, that such $x^{\prime}$ exists by Claim 16 and $\operatorname{agr}(x^{\prime},y^{\prime})=0$, because for all $x^{\prime}$ with $\operatorname{agr}(x^{\prime},y^{\prime})\neq 0$ we have ${\mathcal{F}_{1}^{\prime\prime}}^{\pi^{\prime}}(R^{\prime}\setminus H_{1}\rightarrow x^{\prime})=\emptyset$. Let $\mathcal{F}^{\prime\prime\prime}_{1}={\mathcal{F}_{1}^{\prime\prime}}^{\pi^{\prime}}(R^{\prime}\setminus H_{1}\rightarrow x^{\prime})$. By Claim 5 we get
$\nu^{\prime\prime}_{[n]\setminus(R\cup R^{\prime}\cup H_{1})}(\mathcal{F}^{\prime\prime\prime}_{1})=\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup R^{\prime}\cup H_{1})}(\mathcal{F}^{\prime\prime\prime}_{1})\geqslant\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup H_{1})}({\mathcal{F}_{1}^{\prime\prime}}^{\pi^{\prime}})\geqslant\nu^{\prime}_{[n]\setminus(R\cup H_{1})}(\mathcal{F}^{\prime\prime}_{1})\geqslant\frac{1}{4}.$
@@ 5 Case $m\geqslant\operatorname{poly}(t)$, $n\geqslant\operatorname{poly}(t)\log m$
We need to check that conditions of Corollary 32 imply the conditions of Theorem 28. The case $t=1$, that is, the case of intersecting family of codes, is considered in [30]. Thus, we may assume $t\geqslant 2$. Requirements (6), (7), (8) from Theorem 28 have a form $m\geqslant\text{poly}(t,\log m)$ and the polynomial from the requirement (7) has the highest degree $2C_{2}(2)<300$ in $t$. Thus, there exists a polynomial $m_{0}(t)$ of degree at most $300$ such that $m\geqslant m_{0}(t)$ implies requirements (6), (7), (8). Finally, for $m>e$ the condition (5) is a consequence of a condition $n>n_{0}(t)\ln(m)$, where
@@ Appendix A Proof of Lemma 12
Lemma 12 is a quantitative analogue of Theorem 2.3 from [42] and may be obtained by a verbatim repetition of its proof. We will prove the lemma by contradiction in the following way:
@@
One part of the transformation in Step 1 will be the following compression operator, defined in Section 4.2 of [42].
@@
Definition 9 (Definition 4.2 from [42]).
@@
The transformation at Step 1 will have the following properties, that are direct consequences of Fact 4.1 and Fact 4.3 from [42]
@@
For Step 2, we will need the following lemma of Ellis, Keller, and Lifshitz [13] to obtain a lower bound on $\mu_{1/2}$ from $\mu_{1/m}$.
@@
Lemma 46 (Fact 4.11 from [42]).